# Multimessenger Astronomy Beyond the Standard Model: New Window from Quantum Sensors

### v1

@@ I Introduction
A diverse range of transient astrophysical sources can emit bursts of ULB fields (see e.g. Eby:2024mhd for discussion). This includes explosive emission from boson stars (bosenovae) Eby:2016cnq ; Levkov:2016rkk ; Eby:2021ece ; Arakawa:2023gyq ; Arakawa:2024lqr , superradiant emission from rotating binary black holes Yoshino:2012kn ; Baumann:2021fkf ; Takahashi:2024fyq and supernovae Raffelt:1999tx , among others. These scenarios offer opportunities to detect new quantum fields that may not constitute a significant fraction of DM. The direct detection of ULB fields depends on their local density near the experiment. If the ambient ULB density is significantly lower than that of DM, conventional direct detection experiments will have suppressed signals. However, astrophysical transient events provide a promising avenue for enhanced detection prospects, as they could lead to a temporary increase in the local ULB density at terrestrial or space-based experimental sites. With numerous scenarios predicting ULB production, further studies and numerical simulations are required to precisely characterize their spectral emission properties and event rates.
@@
Importantly, ULBs can experience significant screening effects during propagation, severely limiting their detectability in multimessenger astronomy. This is particularly relevant for quadratic ULB couplings, which induce a back-action effect from SM fields, affecting both ULB propagation and detection Stadnik:2021qyf . Screening effects can also play a prominent role in cosmology and dark energy theories with scalar fields Brax:2013ida , including chameleons Khoury:2003rn , K-mouflage Babichev:2009ee ; Brax:2014wla , dilatons Brax:2010gi , and symmetrons Hinterbichler:2011ca . These effects have been the subject of various experimental searches with laboratory experiments (e.g. Jaffe:2016fsh ; yin:2025 ) and have also been investigated in the context of DM Derevianko:2013oaa ; Hees:2018fpg . Screening mechanisms are critical in determining the feasibility of multimessenger astronomy, as in high-density environments, back-action effects can lead to severe screening, causing significant delays between astrophysical messengers and ULB signals—potentially that can exceed the live-time of experiments. In extreme cases, exponential suppression due to screening could render ULB detection entirely unfeasible.
@@ III.1 Case with $\beta>0$
$d_{i,{\rm crit}}^{(2)}$ $=M_{\text{pl}}^{2}\dfrac{\mathcal{R}^{-2}+\omega^{2}-m_{\phi}^{2}}{8\pi\rho_{i}}$ (21)
$\dfrac{1}{f_{a,{\rm crit}}^{2}}$ $=\dfrac{\mathcal{R}^{-2}+\omega^{2}-m_{a}^{2}}{\dfrac{c_{\gamma}\rho_{\gamma}\alpha}{2\pi^{2}}+c_{e}\rho_{e}}~{}.$ (22)
For the Earth,we approximate the density to be a constant $\rho_{E}=M_{E}/V_{E}\simeq 5.5~{}{\rm g/cm^{3}}$. However, we note that densities can be significantly higher in the Earth’s core, and our analysis can be further refined by incorporating a full Earth structure model Dziewonski:1981xy . We then obtain estimates for the critical coupling values considering energies of $\omega\lesssim 1/\mathcal{R}_{\oplus}^{2}$ as $d_{e,{\rm crit}}^{(2)\,\oplus}\simeq 2.2\times 10^{13}$, $d_{m_{e},{\rm crit}}^{(2)\,\oplus}\simeq 1.7\times 10^{14}$, $d_{g,{\rm crit}}^{(2)\,\oplus}\simeq 4\times 10^{10}$ and $1/f_{a,{\rm crit}}\simeq 2.5\times 10^{-13}$ GeV-1, assuming conservative values of $c_{\gamma},c_{e}=1$. Notably, this value of $1/f_{a,{\rm crit}}$ is already excluded by constraints from observations of the Sun and white dwarfs Hook:2017psm ; DiLuzio:2021pxd .
@@ IV Source Emission
$\begin{cases}t_{*}\delta\omega\gtrsim 1,&\\ t_{*}\omega\gtrsim 2\pi~{}.&\\ \end{cases}$ (32)
To determine the feasibility of multi-messenger astronomy under optimal conditions, we consider the best-case scenario. The momentum spread influences wave spreading—causing additional signal suppression proportional to $\sim 1/R$ in addition to the $\sim 1/R^{2}$ from isotropic signal emission. Hence, for our analysis we assume minimal uncertainty of $\delta\omega=1/t_{*}$.
@@ V.3 Wave Spreading
Figure 5: Duration of the signal at the detector, $\tilde{t}_{*}$, as a function of the time delay, $\delta t$. Different benchmarks for energy and intrinsic burst duration are shown, distinguished by colors and linestyles, respectively. The gray shaded region is excluded by the uncertainty principle relating the intrinsic burst duration and the energy of the emitted ULBs. All curves assume the minimal uncertainty case, $\delta\omega=1/t_{*}$. For larger $\delta\omega$, the nonzero-slope portions of the curves would shift upward linearly.
@@
In Fig. 5, we show the signal duration at the detector, $\tilde{t}_{*}$, as a function of the time delay, $\delta t$, for energy and intrinsic burst time benchmarks of $\omega={10^{-17},10^{-14},10^{-11}}$ eV and $t_{*}=\{1,100\}$ s, respectively. As before, we assume the minimal uncertainty case, $\delta\omega=1/t_{*}$. As illustrated, a larger $\delta\omega$ would result in increased wave spreading and, consequently, a longer signal duration for a given time delay. As a reference, for time delays that allow correlations with standard astrophysical messenger signals, we consider that the signal duration must satisfy $\tilde{t}_{*}\lesssim\mathrm{day}$.
@@ VI Detection Prospects
assuming minimal uncertainty, $\delta\omega=1/t_{*}$. This wave packet has a spatial extent $\tilde{x}_{*}=\tilde{t}_{*}$, an average energy $\omega$, and an energy spread of $\delta\omega$. Consequently, as the signal passes the detector, it exhibits a chirp with a time-dependent frequency. The detection of a signal depends on the signal duration $\tilde{t}_{*}$, the scalar field density at the detector $\rho_{*}$, the detector’s sensitivity as a function of frequency, the coupling type and strength, and the integration time. In Fig. 6 we illustrate a wavepacket described by Eq. (VI) for two different distances from source, $R_{1}$ and $R_{2}$, corresponding to propagation times, $t_{\phi}^{\prime}(R_{1})$ and $t_{\phi}^{\prime}(R_{2})$, respectively. The attenuated amplitude, chirped frequency and enhanced spatial extent of the wave packet are demonstrated.
@@
Here, we discuss representative benchmark scenarios to illustrate the potential sensitivity reach of proposed experiments. Given the diversity of detector types for the various couplings, instead of exhaustive characterization of all possible experimental setups we provide a broad overview of the parameter space that could be probed, leaving detailed, detector-specific analyses for future work. We assume that the integration time for experiments searching for the short duration bursts is $t_{\rm int,*}=1$ day throughout [^7] Our analyses can be readily applied to other target values of $t_{\rm int}*$..
@@ VI.1 Scalars
where $\tau_{\rm DM}=2\pi/(m_{\phi}v^{2})$ is the coherence time of non-relativistic, ambient DM with virial velocity $v_{\text{vir}}\simeq 10^{-3}$ Cooray:2002dia , and $\tau_{*}$ is the effective coherence time of the relativistic $\phi$ fields that originate from an astrophysical burst event.
@@ VI.2 ALPs
For all couplings, we assume $t_{\mathrm{int,*}}=1~{}\mathrm{day}$ to optimize searches for bursts lasting a day or less, which is what is required for time delays $\delta t$ of less than a day, as per Fig. 5.
@@ VII Multimessenger Astronomy
Figure 7: Sensitivity to quadratic dilatonic scalar coupling to photons, $d_{e}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by direct $\phi$ production from supernova photons olive2008environmental . The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (50). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, given all constraints and benchmarks.
Figure 8: Sensitivity to quadratic dilatonic scalar coupling to photons, $d_{e}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by direct $\phi$ production from supernova photons olive2008environmental . Blue solid, dashed, and dotted lines show critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line indicates the projected sensitivity of future experiments with $\eta=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy limits and the uncertainty principle. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
$d_{e}^{(1)}$ $1\times 10^{-4}$  Berge:2017ovy $\cdots$ $\cdots$ $10^{-8}$ ($10^{-4}$) $10^{-6}$ ($2\times 10^{-2}$) 11 (14)
$d_{m_{e}}^{(1)}$ $1\times 10^{-3}$  Berge:2017ovy $\cdots$ $\cdots$ $8\times 10^{-3}$ ($8\times 10^{1}$) $8\times 10^{-1}$ ($1\times 10^{4}$) 12 (15)
$d_{g}^{(1)}$ $3\times 10^{-6}$  Berge:2017ovy $\cdots$ $\cdots$ $8\times 10^{-10}$ ($8\times 10^{-6}$) $8\times 10^{-8}$ ($10^{-3}$) 13 (16)
$d_{e}^{(2)}$ $1\times 10^{30}$  olive2008environmental $2\times 10^{13}$ $6\times 10^{27}$ $2\times 10^{7}$ ($2\times 10^{15}$) $2\times 10^{11}$ ($2\times 10^{19}$) 7 (8)
$d_{m_{e}}^{(2)}$ $5\times 10^{31}$  olive2008environmental $2\times 10^{14}$ $4\times 10^{28}$ $10^{13}$ ($10^{21}$) $10^{17}$ ($9\times 10^{24}$) 17 (19)
$d_{g}^{(2)}$ $5\times 10^{28}$  olive2008environmental $4\times 10^{10}$ $10^{25}$ $10^{6}$ ($10^{14}$) $10^{10}$ ($9\times 10^{17}$) 18 (20)
@@
$g_{\phi\gamma\gamma}\ \rm(GeV^{-1})$ $1\times 10^{-13}$  sisk2022new $\cdots$ $\cdots$ $9\times 10^{-13}$ ($6\times 10^{-11}$) $6\times 10^{-13}$ ($3\times 10^{-12}$) 22 (26)
$g_{\phi e}\ \rm(GeV^{-1})$ $1\times 10^{-10}$  capozzi2020axion $\cdots$ $\cdots$ $2\times 10^{-9}$ ($1\times 10^{-7}$) $1\times 10^{-9}$ ($8\times 10^{-9}$) 21 (25)
$g_{\phi PP}\ \rm(GeV^{-1})$ $6\times 10^{-10}$  lella2024getting $\cdots$ $\cdots$ $7\times 10^{-14}$ ($7\times 10^{-12}$) $7\times 10^{-14}$ ($3\times 10^{-12}$) 23 (27)
$g_{\phi NN}\ \rm(GeV^{-1})$ $1\times 10^{-9}$  buschmann2022upper $\cdots$ $\cdots$ $6\times 10^{-12}$ ($6\times 10^{-10}$) $6\times 10^{-12}$ ($1\times 10^{-10}$) 9 (10)
$g_{\phi P\gamma}\ \rm(GeV^{-2})$ $7\times 10^{-9}$  caloni2022novel $\cdots$ $\cdots$ $2\times 10^{-15}$ ($2\times 10^{-13}$) $2\times 10^{-15}$ ($2\times 10^{-14}$) 24 (28)
Table 1: Sensitivity to SM-$\phi$ couplings from transient astrophysical events from quantum sensors, considering reference Galactic source at $R=10$ kpc with $E_{\rm tot}=10^{-2}M_{\odot}$ and extragalactic source at $R=10$ Mpc with $E_{\rm tot}=M_{\odot}$. Strongest existing constraints as well as critical coupling values for Earth ($d_{i,\ \rm crit}^{\oplus}$) and apparatus ($d_{i,\ \rm crit}^{\rm app}$) are displayed. For scalars, we consider $m_{\phi}=10^{-21}$ eV and $m_{\phi}=10^{-18}$ eV for both sources. For axions, $m_{\phi}=10^{-20}$ eV and $m_{\phi}=10^{-15}$ eV, and $m_{\phi}=10^{-15}$ eV and $m_{\phi}=10^{-10}$ eV, for the two sources respectively. The couplings shaded green are discussed in Sec. VII, while the others are discussed in App. B. The values in parentheses correspond to the extragalactic source benchmark. For a given coupling row, when the sensitivity reach is lower than the strongest constraint, there exists viable observation parameter space for our benchmark choices and assumptions. The dots $\cdots$’ represent that this is not applicable for a given coupling.
@@
Projected experimental sensitivities are shown as black lines, rescaled from ultralight experimental benchmarks using Eqs. (47), (48), (VI.2), and (VI.2). Note that due to the time factors in Eqs. (46), (VI.2), and (VI.2), the sensitivity projections have discontinuities in their slopes. This is associated with transitions between different timescale regimes, such as $\tilde{t}_{*}>t_{\rm int,*}$ or $\tilde{t}_{*}<t_{\rm int,*}$. Gray regions reflect astrophysical or equivalence principle (EP) constraints. Green regions in the plots represent the parameter space where multimessenger astronomy is feasible, taking into account all constraints, physical conditions, and screening effects, based on the chosen benchmark parameters and the best projected sensitivities of future experiments.

### v2

@@ I Introduction
A diverse range of transient astrophysical sources can emit bursts of ULB fields (see e.g. Eby:2024mhd for discussion). This includes explosive emission from boson stars (bosenovae) Eby:2016cnq ; Levkov:2016rkk ; Eby:2021ece ; Arakawa:2023gyq ; Arakawa:2024lqr , superradiant emission from rotating binary black holes Yoshino:2012kn ; Baumann:2021fkf ; Takahashi:2024fyq and supernovae Raffelt:1999tx ; deGiorgi:2024pjb , among others. These scenarios offer opportunities to detect new quantum fields that may not constitute a significant fraction of DM. The direct detection of ULB fields depends on their local density near the experiment. If the ambient ULB density is significantly lower than that of DM, conventional direct detection experiments will have suppressed signals. However, astrophysical transient events provide a promising avenue for enhanced detection prospects, as they could lead to a temporary increase in the local ULB density at terrestrial or space-based experimental sites. With numerous scenarios predicting ULB production, further studies and numerical simulations are required to precisely characterize their spectral emission properties and event rates.
@@
Importantly, ULBs can experience significant screening effects during propagation, severely limiting their detectability in multimessenger astronomy. This is particularly relevant for quadratic ULB couplings, which induce a back-action effect from SM fields, affecting both ULB propagation and detection Stadnik:2021qyf . Screening effects can also play a prominent role in cosmology and dark energy theories with scalar fields Brax:2013ida , including chameleons Khoury:2003rn , K-mouflage Babichev:2009ee ; Brax:2014wla , dilatons Brax:2010gi , and symmetrons Hinterbichler:2011ca . These effects have been the subject of various experimental searches with laboratory experiments (e.g. Jaffe:2016fsh ; yin:2025 ) and have also been investigated in the context of DM Derevianko:2013oaa ; Hees:2018fpg , including the behavior of quadratically-coupled DM around compact objects like the Earth burrage2024time ; bauer2024validity ; del2025enhanced ; banerjee2025momentum .
Screening mechanisms are critical in determining the feasibility of multimessenger astronomy, as in high-density environments, back-action effects can lead to severe screening, causing significant delays between astrophysical messengers and ULB signals—potentially that can exceed the live-time of experiments. In extreme cases, exponential suppression due to screening could render ULB detection entirely unfeasible.
@@ III.1 Case with $\beta>0$
$d_{i,{\rm crit}}^{(2)}$ $=M_{\text{pl}}^{2}\dfrac{(2\mathcal{R})^{-2}+\omega^{2}-m_{\phi}^{2}}{8\pi\rho_{i}}$ (21)
$\dfrac{1}{f_{a,{\rm crit}}^{2}}$ $=\dfrac{(2\mathcal{R})^{-2}+\omega^{2}-m_{a}^{2}}{\dfrac{c_{\gamma}\rho_{\gamma}\alpha}{2\pi^{2}}+c_{e}\rho_{e}}~{}.$ (22)
For the Earth,we approximate the density to be a constant $\rho_{E}=M_{E}/V_{E}\simeq 5.5~{}{\rm g/cm^{3}}$. However, we note that densities can be significantly higher in the Earth’s core, and our analysis can be further refined by incorporating a full Earth structure model Dziewonski:1981xy . We then obtain estimates for the critical coupling values considering energies of $\omega\lesssim 1/(2\mathcal{R}_{\oplus})^{2}$ as $d_{e,{\rm crit}}^{(2)\,\oplus}\simeq 3\times 10^{10}$, $d_{m_{e},{\rm crit}}^{(2)\,\oplus}\simeq 2\times 10^{11}$, $d_{g,{\rm crit}}^{(2)\,\oplus}\simeq 6\times 10^{7}$ and $1/f_{a,{\rm crit}}\simeq 6\times 10^{-14}$ GeV-1, assuming conservative values of $c_{\gamma},c_{e}=1$. Notably, this value of $1/f_{a,{\rm crit}}$ is already excluded by constraints from observations of the Sun, white dwarfs, and neutron stars balkin2024white ; gomez2024constraining ; DiLuzio:2021pxd .
@@ IV Source Emission
$\begin{cases}t_{*}\delta\omega\gtrsim 1/2,&\\ t_{*}\omega\gtrsim 2\pi~{}.&\\ \end{cases}$ (32)
To determine the feasibility of multi-messenger astronomy under optimal conditions, we consider the best-case scenario. The momentum spread influences wave spreading—causing additional signal suppression proportional to $\sim 1/R$ in addition to the $\sim 1/R^{2}$ from isotropic signal emission. Hence, for our analysis we assume minimal uncertainty of $\delta\omega=1/(2t_{*})$.
@@ V.3 Wave Spreading
Figure 5: Duration of the signal at the detector, $\tilde{t}_{*}$, as a function of the time delay, $\delta t$. Different benchmarks for energy and intrinsic burst duration are shown, distinguished by colors and linestyles, respectively. The gray shaded region is excluded by the uncertainty principle relating the intrinsic burst duration and the energy of the emitted ULBs. All curves assume the minimal uncertainty case, $\delta\omega=1/(2t_{*})$. For larger $\delta\omega$, the nonzero-slope portions of the curves would shift upward linearly.
@@
In Fig. 5, we show the signal duration at the detector, $\tilde{t}_{*}$, as a function of the time delay, $\delta t$, for energy and intrinsic burst time benchmarks of $\omega={10^{-17},10^{-14},10^{-11}}$ eV and $t_{*}=\{1,100\}$ s, respectively. As before, we assume the minimal uncertainty case, $\delta\omega=1/(2t_{*})$. As illustrated, a larger $\delta\omega$ would result in increased wave spreading and, consequently, a longer signal duration for a given time delay. As a reference, for time delays that allow correlations with standard astrophysical messenger signals, we consider that the signal duration must satisfy $\tilde{t}_{*}\lesssim\mathrm{day}$.
@@ VI Detection Prospects
assuming minimal uncertainty, $\delta\omega=1/(2t_{*})$. This wave packet has a spatial extent $\tilde{x}_{*}=\tilde{t}_{*}$, an average energy $\omega$, and an energy spread of $\delta\omega$. Consequently, as the signal passes the detector, it exhibits a chirp with a time-dependent frequency. The detection of a signal depends on the signal duration $\tilde{t}_{*}$, the scalar field density at the detector $\rho_{*}$, the detector’s sensitivity as a function of frequency, the coupling type and strength, and the integration time. In Fig. 6 we illustrate a wavepacket described by Eq. (VI) for two different distances from source, $R_{1}$ and $R_{2}$, corresponding to propagation times, $t_{\phi}^{\prime}(R_{1})$ and $t_{\phi}^{\prime}(R_{2})$, respectively. The attenuated amplitude, chirped frequency and enhanced spatial extent of the wave packet are demonstrated.
@@
Here, we discuss representative benchmark scenarios to illustrate the potential sensitivity reach of proposed experiments. Given the diversity of detector types for the various couplings, instead of exhaustive characterization of all possible experimental setups we provide a broad overview of the parameter space that could be probed, leaving detailed, detector-specific analyses for future work. We assume that the integration livetime for experiments searching for the short duration ULB transient bursts is $t_{\rm int,*}=1$ day throughout, although our analyses can be readily applied to other target values of $t_{\rm int}*$.
@@ VI.1 Scalars
where $\tau_{\rm DM}=2\pi/(m_{\rm DM}v^{2})$ is the coherence time of non-relativistic, ambient DM with virial velocity $v_{\text{vir}}\simeq 10^{-3}$ Cooray:2002dia , and $\tau_{*}$ is the effective coherence time of the relativistic $\phi$ fields that originate from an astrophysical burst event. In our analysis we compare the sensitivity of conventional DM searches with that of ULB transient searches at a common experimental frequency $f$, where the $\phi$-field energy is $\omega=2\pi f$. As we scan over frequency we set corresponding non-relativistic DM mass $m_{\rm DM}=\omega$. Hence, $m_{\rm DM}$ is distinct from benchmark ULB mass $m_{\phi}$. Owing to this distinction, ULB transient searches can be intrinsically more sensitive to lower $m_{\phi}$ values within any fixed frequency band.
@@ VI.2 ALPs
For all couplings, we assume $t_{\mathrm{int,*}}=1~{}\mathrm{day}$ to optimize searches for bursts lasting a day or less, which is what is required for time delays $\delta t$ of less than a day, as per Fig. 5. However, for different signal types and durations, one could choose a different $t_{\rm int,*}$, changing the experimental sensitivities according to Eqs. (46), (VI.2), and Eq. (VI.2).
@@ VII Multimessenger Astronomy
Figure 7: Sensitivity to quadratic dilatonic scalar coupling to photons, $d_{e}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by direct $\phi$ production from supernova photons olive2008environmental . The blue solid, dashed, and dotted lines indicate critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line shows the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (50). The black hatched and brown regions are excluded due to the scalar energy being less than its mass and the uncertainty principle, respectively. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, given all constraints and benchmarks.
Figure 8: Sensitivity to quadratic dilatonic scalar coupling to photons, $d_{e}^{(2)}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{Mpc}$ from Earth, releasing a total energy of $E_{\rm total}=M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-21}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-18}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by direct $\phi$ production from supernova photons olive2008environmental . Blue solid, dashed, and dotted lines show critical screening from the Earth, atmosphere, and experimental apparatus, respectively. The black line indicates the projected sensitivity of future experiments with $\eta_{\mathrm{DM}}=10^{-19}/5900$, based on a Thorium isomer nuclear clock proposal beeks2024fine , scaled using Eq. (50). The black hatched and brown regions are excluded by scalar energy limits and the uncertainty principle. Dashed purple and red lines represent detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.
@@
$d_{e}^{(1)}$ $1\times 10^{-4}$  Berge:2017ovy $\cdots$ $\cdots$ $2\times 10^{-6}$ ($3\times 10^{-4}$) $10^{-4}$ ($2\times 10^{-1}$) 11 (14)
$d_{m_{e}}^{(1)}$ $1\times 10^{-3}$  Berge:2017ovy $\cdots$ $\cdots$ $1.1$ ($2\times 10^{2}$) $6\times 10^{1}$ ($1\times 10^{5}$) 12 (15)
$d_{g}^{(1)}$ $3\times 10^{-6}$  Berge:2017ovy $\cdots$ $\cdots$ $1\times 10^{-7}$ ($2\times 10^{-5}$) $6\times 10^{-6}$ ($1\times 10^{-2}$) 13 (16)
$d_{e}^{(2)}$ $1\times 10^{31}$  olive2008environmental $3\times 10^{10}$ $1\times 10^{24}$ $5\times 10^{9}$ ($8\times 10^{14}$) $4\times 10^{12}$ ($4\times 10^{18}$) 7 (8)
$d_{m_{e}}^{(2)}$ $5\times 10^{31}$  olive2008environmental $2\times 10^{11}$ $1\times 10^{25}$ $3\times 10^{15}$ ($5\times 10^{20}$) $2\times 10^{18}$ ($2\times 10^{24}$) 17 (19)
$d_{g}^{(2)}$ $5\times 10^{28}$  olive2008environmental $6\times 10^{7}$ $2\times 10^{21}$ $3\times 10^{8}$ ($5\times 10^{13}$) $2\times 10^{11}$ ($2\times 10^{17}$) 18 (20)
@@
$g_{\phi\gamma\gamma}\ \rm(GeV^{-1})$ $1\times 10^{-13}$  sisk2022new $\cdots$ $\cdots$ $2\times 10^{-15}$ ($3\times 10^{-13}$) $2\times 10^{-15}$ ($3\times 10^{-13}$) 22 (26)
$g_{\phi e}\ \rm(GeV^{-1})$ $1\times 10^{-10}$  capozzi2020axion $\cdots$ $\cdots$ $5\times 10^{-13}$ ($8\times 10^{-11}$) $5\times 10^{-13}$ ($8\times 10^{-11}$) 21 (25)
$g_{\phi PP}\ \rm(GeV^{-1})$ $6\times 10^{-10}$  lella2024getting $\cdots$ $\cdots$ $2\times 10^{-15}$ ($3\times 10^{-13}$) $2\times 10^{-15}$ ($3\times 10^{-13}$) 23 (27)
$g_{\phi NN}\ \rm(GeV^{-1})$ $1\times 10^{-9}$  buschmann2022upper $\cdots$ $\cdots$ $5\times 10^{-14}$ ($9\times 10^{-12}$) $5\times 10^{-14}$ ($9\times 10^{-12}$) 9 (10)
$g_{\phi P\gamma}\ \rm(GeV^{-2})$ $7\times 10^{-9}$  caloni2022novel $\cdots$ $\cdots$ $1\times 10^{-17}$ ($2\times 10^{-15}$) $1\times 10^{-17}$ ($2\times 10^{-15}$) 24 (28)
Table 1: Sensitivity to SM-$\phi$ couplings from transient astrophysical events from quantum sensors, considering reference Galactic source at $R=10$ kpc with $E_{\rm tot}=10^{-2}M_{\odot}$ and extragalactic source at $R=10$ Mpc with $E_{\rm tot}=M_{\odot}$. The sensitivities displayed meet the time delay and emission constraints (defined by the minimum of the green regions in the Figures in the final column). Strongest existing constraints as well as critical coupling values for Earth ($d_{i,\ \rm crit}^{\oplus}$) and apparatus ($d_{i,\ \rm crit}^{\rm app}$) are displayed. For scalars, we consider $m_{\phi}=10^{-21}$ eV and $m_{\phi}=10^{-18}$ eV for both sources. For axions, $m_{\phi}=10^{-20}$ eV and $m_{\phi}=10^{-15}$ eV, and $m_{\phi}=10^{-15}$ eV and $m_{\phi}=10^{-10}$ eV, for the two sources respectively. For all couplings shown, we consider a benchmark experimental integrated livetime of $t_{\rm int,*}=1$ day. The couplings shaded green are discussed in Sec. VII, while the others are discussed in App. B. The values in parentheses correspond to the extragalactic source benchmark. For a given coupling row, when the sensitivity reach is lower than the strongest constraint, there exists viable observation parameter space for our benchmark choices and assumptions. The dots $\cdots$’ indicate that a critical value is not applicable for the given coupling.
@@
Projected experimental sensitivities are shown as black lines, rescaled from ultralight experimental benchmarks using Eqs. (47), (48), (VI.2), and (VI.2). Note that due to the time factors in Eqs. (46), (VI.2), and (VI.2), the sensitivity projections have discontinuities in their slopes. This is associated with transitions between different timescale regimes, such as $\tilde{t}_{*}>t_{\rm int,*}$ or $\tilde{t}_{*}<t_{\rm int,*}$, as well as $\tau_{\rm DM}>t_{\rm int,DM}$ or $\tau_{\rm DM}<t_{\rm int,DM}$. Gray regions reflect astrophysical or equivalence principle (EP) constraints. Green regions in the plots represent the parameter space where multimessenger astronomy is feasible, taking into account all constraints, physical conditions, and screening effects, based on the chosen benchmark parameters and the best projected sensitivities of future experiments.